OpenStudy (anonymous):

Solve for the variable 3w = 4 1/5 ? The four is a whole number 1/5 is a fraction.

OpenStudy (anonymous):

To simplify the problem i would first convert all the whole numbers to fractions of 5ths.

OpenStudy (cgreenwade2000):

3w =/= 4/5

OpenStudy (anonymous):



OpenStudy (anonymous):

You mulitply 4 and 1 first correct , and then five and so on?

OpenStudy (jwheele1):

is this your equation? 3w = 4 * (1/5)

OpenStudy (cgreenwade2000):

No. It doesn't say four times one-fifth. It says four and one fifth

OpenStudy (anonymous):

o i see...

OpenStudy (cgreenwade2000):

Which means you need to turn the four into a fraction with a denominator of 5 and add it to 1/5

OpenStudy (jwheele1):

\[3w=4\frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

Yes that is the equation , I got 7/5 ? Is that correct?

OpenStudy (cgreenwade2000):

Yes Gryffin.

OpenStudy (anonymous):

Thank You. c:

OpenStudy (cgreenwade2000):

Just remember how fractions and mixed numbers work. There are websites that can help you build your skills.

OpenStudy (jwheele1):

Well the first thing I'd do is take that ugly mixed number and make it a regular decimal or an improper fraction or something You take the whole number and multiply it by your denominator; so 4*5 = 20. Than you add the numerator to your number so 20 + 1 = 21. 21 becomes your new numerator and the denominator stay the orignal 5. You new fraction is 21/5. Your new equation is 3w = (20/5)

OpenStudy (cgreenwade2000):

21/5 jwheele

OpenStudy (jwheele1):

This is true, lol

OpenStudy (jwheele1):

well...you guys got the answer anyways. lol